OpenStudy (anonymous):

A motorcycle is capable of accelerating at 1.5 m/s2. Starting from rest, how far can it travel in 0.5 seconds? A. 0.38 m B. 0.44 m C. 0.19 m D. 0.66 m

OpenStudy (aaronq):

Use the kinematic equations of motion http://hyperphysics.phy-astr.gsu.edu/hbase/mot.html

OpenStudy (michele_laino):

here we have to apply this formula: \[\Large s = \frac{1}{2}a{t^2} = \frac{1}{2} \times 1.5 \times {0.5^2} = ...meters\] where s is the requested traveled space by the motorcycle